OpenStudy (anonymous):

Please, Please help. Simplify each expression by rationalizing the denomination (2-√(3)) / (√(2) + √(6)). Show steps please.

OpenStudy (anonymous):

multiplied by \[(\sqrt{2}-\sqrt{6})/\]

OpenStudy (anonymous):

sorry typo, \[(\sqrt{2}-\sqrt{6})/(\sqrt{2}-\sqrt{6}) \]

OpenStudy (anonymous):

the result will be \[(2\sqrt{2} - 2\sqrt{6} -\sqrt{6} + \sqrt{18})/2-6\] \[= (5\sqrt{2}-3\sqrt{6})/(-4) = 3\sqrt{6}-5\sqrt{2}/4\]

OpenStudy (anonymous):

\[(a \pm b)/(\sqrt{c}+\sqrt{d}) \] to razionalize it, multiply it with \[(\sqrt{c}-\sqrt{d})/(\sqrt{c}-\sqrt{d})\] and if you have \[(a \pm b)/(\sqrt{c}-\sqrt{d}) \] to razionalize it, multiply it with \[(\sqrt{c}+\sqrt{d})/(\sqrt{c}+\sqrt{d})\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

here your answer